OpenStudy (anonymous):

simplify as far as possible: (2x^2 +x -12) / (x^2 - 16)....i'm not sure how to simplify the top

OpenStudy (anonymous):

It doesnt look like it will simplify. Could that be a 2x at the top?

OpenStudy (anonymous):

its (2x)^2 + x - 12

OpenStudy (anonymous):

i thought maybe i can complete the square

OpenStudy (anonymous):

not much point in completing the square when there's a denominator. is it (2x)^2 or 2x^2?

OpenStudy (anonymous):

2x^2

OpenStudy (anonymous):

\[(2x ^{2}+ x - 12 ) / (x ^{2} -16)\]

OpenStudy (anonymous):

tahts wat it looks like

OpenStudy (anonymous):

ok. well come to think of it that doesnt make much of a difference. I dont think it can be simplified further if your teacher just wants you to factor top and bottom and cancel like terms

OpenStudy (anonymous):

i see....can u help me with anotehr simplify problem

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

\[x ^{3/2} ( x+ x ^{5/2} - x ^{2})\]

OpenStudy (anonymous):

ok, so basically we are just going to do some distribution when you multiply terms with exponents, you are really just adding the exponents for example 3^3 = 3 * 3 *3 So 3^3 * 3^4 = 3*3*3*3*3*3*3 = 3^7 this works with fractions too! so 3/2 +1 = 5/2, 5/2 + 3/2= 4, and 2+3/2 = 7/2 so simplified fully it should be x^(5/2) + x^4 - x^(7/2)

OpenStudy (anonymous):

Im sorry it takes so long but Im a poor typist

OpenStudy (anonymous):

its okay...you are the one helping :)

OpenStudy (anonymous):

i keep asking this q but i kee[ getting differnt answers. what is \[\sqrt{x} / x\]

OpenStudy (anonymous):

teachers dont ask for exponents to be in fractions, so it would just be 2x^2+x-12/(x+4)(x-4)

OpenStudy (anonymous):

you're a bit late ivy

OpenStudy (anonymous):

alright im tired of typing. express sqrt(x) as x^something

OpenStudy (anonymous):

lets do this interactively

OpenStudy (anonymous):

dangg man... way to point me out! haha i was just saying